OpenStudy (mathsucks321):

What is the solution for the following equation? 6(x + 5) + 2 = 8(2x - 5) + 22 A. 4.8 B. ­5 C. 3 D. 5

OpenStudy (mathsucks321):

Just have to find x right?

OpenStudy (3mar):

Yes, and I am sure you can!

OpenStudy (mathsucks321):

I know the answer then

OpenStudy (mathsucks321):

x=5

OpenStudy (mathsucks321):

But I have to show how to get it

OpenStudy (3mar):

please.... Can you share your works/steps, please?

OpenStudy (mathsucks321):

I substituted 5 for x, I tried them all first tho, but I have to answer differently

OpenStudy (mathsucks321):

Since I got x=5 and that is right, can you help me get a different equation for it?

OpenStudy (3mar):

Look!!! \(6(x + 5) + 2 = 8(2x - 5) + 22\) start with parenthesis: \(6x + 30 + 2 = 16x - 10 + 22\) Simplify: \(6x + 32 = 16x +12\) \(6x + 32-12 = 16x +12-12\) \(6x + 20= 16x \) \(6x + 20-6x= 16x-6x \) \(20=10x \) \(20÷10=10x÷10 \) \(\Large2=x \)

OpenStudy (mathsucks321):

2??

OpenStudy (3mar):

SORRRRRY 6(x+5)+2=8(2x−5)+22 start with parenthesis: 6x+30+2=16x−40+22 Simplify: 6x+32=16x-18 32+18=16x-6x 50=10x \(x=5\)

OpenStudy (mathsucks321):

lol better

OpenStudy (mathsucks321):

Thank you <3

OpenStudy (3mar):

You are welcome!